A Field Poll Survey reported that 62% of registered voters in a state approved of allowing two people of the same gender to marry and have regular marriage laws apply to them. Among 18 to 39 year olds (registered voters in this state), the approval rating was 78%. Six in ten registered voters in this state said that the upcoming Supreme Court's ruling about the constitutionality of a proposition was either very or somewhat important to them. Out of those registered voters who support same-sex marriage, 71% say the ruling is important to them. In this problem, let the following apply.
• C = registered voters who support same-sex marriage
• B = registered voters who say the Supreme Court's ruling about the constitutionality of the proposition is very or somewhat important to them
• A = registered voters who are 18 to 39 years old.
1. Find P(C)= P(C|A)= P(C AND B
1
Expert's answer
2020-09-17T14:12:10-0400

Solution:

• C = California registered voters who support same-sex marriage.

So p(C)= 0.62 from given data.




B and C is California registered voters who support same-sex marriage and who say 

the Supreme Court’s ruling about the constitutionality of California’s 

Proposition 8 is very or somewhat important to them


Hence P(C AND B) = P(B|C)P(C) = 0.71*0.62= 0.4402


P(C/A):

Among 18 to 39 year olds (registered voters in this state), the approval rating was 78%.


Hence p(C/A)= 0. 78